OpenStudy (anonymous):

Simplify 6r · s · 4rt.

OpenStudy (anonymous):

(6r)(s)(4rt) (6r)(s)(4rt) = 24r²st ---->(-2/3x)(3y)(-2x) (-2/3x)(3y)(-2x) = 4x²y = 0

OpenStudy (♪chibiterasu):

Lol wow

OpenStudy (anonymous):

24r^2st

OpenStudy (anonymous):

Simplifying 6r * s * 4rt = 0 Reorder the terms for easier multiplication: 6 * 4r * s * rt = 0 Multiply 6 * 4 24r * s * rt = 0 Multiply r * s 24rs * rt = 0 Multiply rs * rt 24r2st = 0 Solving 24r2st = 0 Solving for variable 'r'. Move all terms containing r to the left, all other terms to the right. Divide each side by '24'. r2st = 0 Simplifying r2st = 0

OpenStudy (♪chibiterasu):

Nice: http://www.geteasysolution.com/6r*s*4rt=

OpenStudy (♪chibiterasu):

lol

OpenStudy (anonymous):

@Paigey

OpenStudy (anonymous):

Thank you (: